OpenStudy (anonymous):

Suppose a triangle has sides 3, 4, and 5. Which of the following must be true? A. The triangle in question is not a right triangle. B. The triangle in question is a right triangle. C. The triangle in question may or may not be a right triangle.

OpenStudy (comm.dan):

A @SantanaG

OpenStudy (anonymous):

The triangle is a right triangle

OpenStudy (anonymous):

A

OpenStudy (anonymous):

because A^2+B^2=c^2

OpenStudy (anonymous):

So B

OpenStudy (anonymous):

oops my bad B lol

OpenStudy (anonymous):

lol Yeah i was like whhatttt? hahah its all good!

OpenStudy (anonymous):

b is correct. A right triangle has one angle of 90 degrees. a^2 + b^2 = c^2 3^2 + 4^2 = 5^2 9 + 16 = 25 25 = 25

OpenStudy (anonymous):

Thanks @Chooch146 and @jpcoogan :)

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

lol ur welcome:]